OpenStudy (anonymous):

A number was added to each value in an original data set to give a new data set with values eleven, 8, 7, twelve, and 7. What number was added to each value if the original mean was 6? 1 2 3 6

OpenStudy (anonymous):

Okay, so first find the mean of the new data.

OpenStudy (anonymous):

11+8+7+12+7 = 45 / 5 = 9

OpenStudy (anonymous):

so your new mean is 9, and your old mean is 6 ... how much did the mean increase?

OpenStudy (anonymous):

by 3

OpenStudy (anonymous):

There ya go :D

OpenStudy (anonymous):

Thanks so much! :D

OpenStudy (anonymous):

No problem!